BEAST9898:

What is the mass defect of lithium-7? Assume the following: Atomic number of lithium = 3 Atomic mass of lithium = 7.016003 atomic mass units. Mass of 1 proton = 1.007276 atomic mass units. Mass of 1 neutron = 1.008665 atomic mass units.

umm:

Welcome to QuestionCove. Any ideas/guesses?

sillybilly123:

\({Li}\): p = 3, n = 4 \(\Delta m = 3 \times 1.007276 + 4 \times 1.008665 - 7.016003 = 0.040485 \)